A 1.0 kg object, A, with a velocity of 4.0m/s to the right, strikes a second object, B, of mass 3.0kg, originally at rest. In the collision, A is deflected from its original direction through an angle of 50°; its speed after the collision is 2.0m/s. Find the angle between B’s velocity after the collision and the original direction of A, and find the speed of B after collision.
1
Expert's answer
2020-07-09T10:33:38-0400
"v-v'\cos{50}=u\cos{\alpha}\\v'\sin{50}=u\sin{\alpha}\\u^2=(v-v'\cos{50})^2+v'^2\sin^2{50}\\u^2=(4-2\cos{50})^2+2^2\sin^2{50}\\u=3.1\frac{m}{s}"

"2\sin{50}=3.1\sin{\alpha}\\\alpha=29\degree"